OpenStudy (anonymous):

Medal always given, help is always appreciated(: How would I solve -sin^2x=2cosx-2 ??

OpenStudy (anonymous):

Hey Velasquez.....do u knw cosx=1-2sin^2x/2

jhonyy9 (jhonyy9):

like a first step sin^2 x +cos^2 x =1 so sin^2 x = 1-cos^2 x and now rewrite sin^2 x =2cosx -2 1 -cos^2 x =2cos x -2 cos^2 x +2cos x -3 =0 than note cos x = t and rewrite it will give you t^2 +2t -3 = 0 so now this you need to solve it for t and wll get like roots t_1 and t_2 what dont forget you need calculing x making cos x = t_1 and cos x =t_2 hope this will help you

OpenStudy (only4u):

ladkiyo ko sirf qn puchna aata hai

OpenStudy (anonymous):

stuck on the last part :/

OpenStudy (anonymous):

quadratic formula

OpenStudy (anonymous):

oh the last part

OpenStudy (jdoe0001):

\(\bf sin^2(x)=2cos(x)-2\\ \quad \\ \textit{recall that }sin^2(\theta)+cos^2(\theta)=1\implies sin^2(\theta)=1-cos^2(\theta)\qquad thus\\ \quad \\ sin^2(x)=2cos(x)-2\implies 1-cos^2(x)=2cos(x)-2\\ \quad \\ 0=cos^2(x)+2cos(x)-1\) factor the quadratic like you'd any other, to get the roots and thus the values for cos(x)

OpenStudy (anonymous):

thanks

jhonyy9 (jhonyy9):

@jdoe0001 how you have got this 0=cos2(x)+2cos(x)−1 ? so because 1 on the right side will be minus 1 and with -2 will add -3 yes ?